Without using table solve the following; cot 75°
Tan 90°
Cosec 15°
 
 
 
 
 
 
 
 
 
 
 
 
 
 

Grade:12th pass

1 Answers

Arun
4 years ago

Cot(-75)=-Cot 75⁰

Cot(A+B)=( CotA CotB-1)/(CotA+CotB)

- Cot 75⁰= -Cot(45⁰+30⁰)

= −[(Cot 45⁰Cot30⁰-1)/(Cot45⁰+Cot30⁰)]

= -(√3–1)/(1+√3)= (1-√3)/(1+√3)

= (1-√3)(1-√3)/(1+√3)(1-√3)

=( 4–2√3)/(-2)

= √3–2